2 charged particles reacted with a force of 0.40 N with a distance of 1.6 cm, the 2nd charge contains 60 nC. Find the value of the first charge.



1
Expert's answer
2022-03-07T12:39:27-0500

"F=\frac{kq_1q_2}{r^2}"

"q_2=\frac{Fr^2}{kq_1}"

Put value


"q_2=\frac{0.40\times0.016^2}{9\times10^9\times60\times10^{-9}}=18.97\times10^{-6}C"